# Leveraging Large Language Models for Active Merchant Non-player Characters

### v2

@@ Abstract
We highlight two significant issues leading to the passivity of current merchant non-player characters (NPCs): pricing and communication. While immersive interactions have been a focus, negotiations between merchant NPCs and players on item prices have not received sufficient attention. First, we define passive pricing as the limited ability of merchants to modify predefined item prices. Second, passive communication means that merchants can only interact with players in a scripted manner. To tackle these issues and create an active merchant NPC, we propose a merchant framework based on large language models (LLMs), called mart, which consists of an appraiser module and a negotiator module. We conducted two experiments to guide game developers in selecting appropriate implementations by comparing different training methods and LLM sizes. Our findings indicate that finetuning methods, such as supervised finetuning (SFT) and knowledge distillation (KD), are effective in using smaller LLMs to implement active merchant NPCs. Additionally, we found three irregular cases arising from the responses of LLMs. We expect our findings to guide developers in using LLMs for developing active merchant NPCs.
Index Terms:
Large language models, Non-playing Characters, Active agents, Negotiation, Role-playing game
I Introduction
@@
Exchanging in-game items is an integral part of open–world role-playing games because the utility of an item often depends on the current character status of a player. For example, a character with high agility may not benefit from an item that enhances agility. Thus, game developers implement an exchange system in their open-world games to enhance gameplay and item utility.
@@
While games commonly feature merchant non-player characters (NPCs) to facilitate item exchanges, these interactions are usually scripted; the merchant only enables players to buy and sell items at fixed prices, but the process differs from real-world transactions. Typically, developers do not allow merchants to alter prices or communicate in free form; instead, they simply present a fixed price to the player without any negotiation. To mirror the real world, We suggest a novel framework that uses large language models (LLMs) and enables merchant NPCs to negotiate with players.
@@
To create a more active merchant NPC, we need to address two issues that cause the merchant to act passively: pricing and communication. First, regarding passive pricing, the merchant has no authority to adjust item prices. Instead, the game developers decide item prices based on the item utility in the game. However, in the real world, sellers can adjust the prices of their goods according to their situation. To the best of our knowledge, gaming industry researchers have not adequately filled this gap. As item descriptions may contain sufficient information about item utility, we use item descriptions to estimate the item’s value. Specifically, we let LLMs appraise game items by observing other game items.
@@
Second, regarding passive communication, merchants can only interact with players in a scripted manner. Researchers have attempted to create more immersive interactions by using LLMs [1, 2]. However, the way merchants communicate has been given limited focus, which remains a one-way interaction. Current merchant NPCs simply display a predefined list of items for players to choose from, and players respond by clicking on items to make a purchase. This purchasing experience is uniform across all interactions with a merchant, regardless of the individual player involved in the transaction. To enable two-way communication, we propose a negotiation style that mirrors real-world transactions. Through this study, we explore whether LLMs can be used to foster a negotiating interaction within the merchant context.
@@
In this paper, we propose a framework for developing a More Active meRchanT NPC, called mart. This framework consists of two main components: appraiser and negotiator. The appraiser module addresses the issue of passive pricing by estimating the value of given items. The negotiator module addresses the passive communication issue by negotiating with players. As game developers have diverse requirements when deploying their games, we conducted experiments to compare potential candidates for each module using a public WoW Classic game item dataset. Specifically, we tested both finetuning methods and $n$-shot prompting methods on the family of Llama 3 models, which come with a wide range of parameter sizes from 1 billion to 405 billion.
This study makes the following contributions:
@@
  • By statistically and qualitatively analyzing the results, we provide several design options for developing active merchants in a game.
II MART Framework
@@
We propose mart [^1] Source code: http://github.com/elu-lab/mart, an LLM-powered framework for active merchant NPC. This framework consists of two main modules: negotiator and appraiser. Inspired by the behavior of real-world merchants, MART negotiates item purchases with players. When a player expresses interest in buying an item, the appraiser module suggests a retail price for the item. Based on this suggestion, the negotiator module begins negotiations with players who wish to buy the item from the merchant NPC. This scenario is different from the purchasing experience of players in current games, as shown in Figure 1.
@@
First, the appraiser module is designed using a language model that interprets an item description to estimate its retail price. The module takes item descriptions presented as natural language sentences, which include details such as required level, effects, and durability. As such descriptions can vary across different games, we opted for a natural language representation to attain versatility. In Section III, we introduce LLM-based appraiser modules using two straightforward approaches and evaluate their performance.
@@
Second, the negotiator module employs a language model to facilitate negotiation dialogues with players. This module receives the item descriptions, retail price, and history of previous conversations. It then generates the appropriate response aimed at persuading the player. Since effective negotiation involves various tactics to persuade players, we experimented LLM-based negotiator modules with two simple approaches that integrate 10 tactics, as detailed in Section IV.
III Appraiser Module
@@
To develop appraisers, we tested two approaches: $n$-shot in-context learning (ICL) and supervised finetuning (SFT). In this section, we outline each approach and illustrate our findings through an experiment on a WoW Classic item dataset.
III-A Two possible approaches
@@
First, to leverage the ability of LLMs to generalize from a set of given examples, we tested the ICL method. ICL involves querying LLMs with relevant demonstrations to improve their reasoning capabilities. Thus, we directly employed LLMs and provided ten random examples in the prompt. We used Llama models, a popular open-source family of LLMs: Llama-3.1 8b, 405b, and 3.2 1b. Note that ICL does not require any additional training to update the LLM parameters.
@@
Second, owing to the substantial computational resources and time required by the 405 billion parameter model, we tested the SFT method on smaller LLMs. To avoid catastrophic forgetting [^2] Catastrophic forgetting indicates a phenomenon that language models forget their parametric knowledge while finetuning them., we froze their parameters and only trained an additional adapter. When we input an item description, the smaller LLM transforms the description into a single latent vector. Thereafter, the additional adapter uses the vector to predict the retail price of the item as a regression problem. We used Llama 3.1 8b and 3.2 1b.
III-B Used Dataset
@@
By extending the WoW Classic item dataset [^3] https://www.kaggle.com/datasets/mylesoneill/classic-world-of-warcraft-auction-data, we created a dataset consisting of a paired list of purchasable items and their prices. Dataset construction followed three steps: First, we extracted 3,376 fundamental items by excluding item derivatives from the crawled dataset to simplify the problem. We then removed 106 items priced below 10 coppers (the cheapest currency unit in WoW) to make enough room for price negotiation during transactions. Finally, we converted all item prices into coppers, using the conversions of 1 gold to 100 silver and 1 silver to 100 coppers. As a result, we collected a total of 3,270 items, with prices ranging from 10 coppers to 57,018 coppers. More than 50% of the items were priced below 1,250 coppers, with a median price of 1,238 coppers and an average price of 3,249 coppers. We further divided the dataset into a training set (80%, 2,616 items), validation set (10%, 327 items), and test set (10%, 327 items).
III-C Evaluation metrics
@@
To evaluate the suitability of the two methods for implementing an appraiser module, we used four metrics: mean absolute percentage error (MAPE), standard deviation, skewness, and unexpected output rate. First, we used MAPE. Due to the wide range of item prices (from 10 to 57,018 coppers), the absolute errors of higher-priced items may overshadow those of lower-priced items. Therefore, we used percentage errors instead of absolute errors. We computed MAPE as the ratio of error to true price, as illustrated in Equation 2, where $\hat{y}_{i}$ is the predicted price and $y^{*}_{i}$ is the true price for an item $i$.
@@
III-D Implementation detail
We implemented four appraisers using the following specifications. First, we used public APIs [^4] http://openrouter.ai to call ICL models. When calling the LLMs, we used default hyperparameter settings, including temperature and top-p values. Second, we implemented the SFT models using PyTorch 2.5.1 [3] and transformers 4.46.2 [4]. As an adapter head for prediction, we applied average pooling on the last hidden state and transformed it into a single prediction using feed-forward layers. Third, we used the early-stopping method for training SFT models. The models are trained with two A6000 GPUs.
III-E Results
TABLE I: Results of Assessing Two Methods of Appraiser
MAPE Standard Skewness Unexpected
Deviation Output Rate
In-context 1b 14.68 44.42 10.23 29.50
Learning 8b 4.34 12.36 6.85 20.49
@@ III-E Results
Fine 1b 3.59 11.57 -5.84 -
Tuning 8b 2.66 11.26 -3.47 -
@@
Table I shows the result of our experiments, comparing ICL and SFT methods. Of the five models, the ICL-405b achieved the best performance. It successfully outputted a retail price (94.8% of the cases) with the lowest MAPE score (1.34%) and produced a few unexpected outputs (5.2%), while it was slightly skewed toward underestimation (-5.06). The second-best model was the SFT-8b. It exhibited a slightly higher MAPE score (2.66%) and the lowest absolute skewness toward underestimation (-3.47). Meanwhile, the SFT-1b demonstrated lower performance than the best model (3.59%) with a similar underestimation (-5.84). These results are different from those of smaller ICL models. The ICL-1b exhibited very high MAPE (14.68%) with highly overestimated prices (10.23) and a high unexpected output rate (29.05%). Similarly, the ICL-8b overestimated prices (6.85) with a moderate level of MAPE (4.34) and a high unexpected output rate (20.59%).
III-F Discussion
@@
We discuss our results in terms of three factors considering game deployment: performance, efficiency, and reliability. First, the appraiser module has the potential to be used in a game because of its performance. Except for ICL-1b, five models demonstrated a MAPE of less than 5%. While the difference may seem substantial for high-priced items, it does not exceed 100 coppers for over half of the items, considering that the median price was 1,238 coppers. As gold coins are more frequently used in the WoW Classic, a difference of 100 coppers is acceptable. Moreover, when using larger ICL models, the appraised prices aligned more closely with the true prices. Therefore, for game developers seeking a more precise appraiser module, ICL-405b is the best option, as it exhibited an appraisal error of less than 16 coppers for the median price.
@@
Second, the SFT method is efficient for developing an appraiser module. Results show that SFT-8b performed much closer to ICL-405b than to ICL-8b; even the SFT-8b and ICL-8b used almost identical number of parameters. Also, the SFT-1b outperformed the two ICL models, ICL-1b and ICL-8b. These results imply that it is possible to use a smaller LLM to implement an appraiser within a low-resource environment by finetuning it. Note that the ICL models achieved high performance by observing ten random pairs of item descriptions and prices. Researchers have reported that the performance of ICL methods can be improved by carefully curating examples or increasing the number of examples[5]. However, this curation demands additional human resources, and increasing the number of examples incurs higher computational costs. By contrast, a dataset for SFT can be generated with significantly fewer resources, using pairs of items and their prices.
@@
Third, it is worth noticing that ICL methods can sometimes be unreliable when playing the role of an appraiser. We observed that ICL models sometimes produce unexpected outputs. For instance, ICL-405b, which employs the largest LLM, generated multiple price candidates or a continuous range of prices, despite our request for a single price output. Moreover, smaller ICL models sometimes totally failed to predict prices. Although we did not adopt any post-processing methods for these outputs, game developers should prevent such failures in their games. Alternatively, developers can use the SFT method, ensuring an LLM predicts a specific price.
IV Negotiator Module
@@
To develop negotiators, we tested two approaches: zero-shot prompting (ZSP) and knowledge distillation (KD). We adopted two methods different from the appraiser module because it is hard to obtain a supervised set of negotiations. In the following subsections, we describe our methods, dataset generation procedure, and details of the experiments. Then, we discuss our findings by comparing these two approaches.
IV-A Two possible approaches

### v3

@@ Abstract
We highlight two significant issues leading to the passivity of current merchant non-player characters (NPCs): pricing and communication. While immersive interactions with active NPCs have been a focus, price negotiations between merchant NPCs and players remain underexplored. First, passive pricing refers to the limited ability of merchants to modify predefined item prices. Second, passive communication means that merchants can only interact with players in a scripted manner. To tackle these issues and create an active merchant NPC, we propose a merchant framework based on large language models (LLMs), called mart, which consists of an appraiser module and a negotiator module. We conducted two experiments to explore various implementation options under different training methods and LLM sizes, considering a range of possible game environments. Our findings indicate that finetuning methods, such as supervised finetuning (SFT) and knowledge distillation (KD), are effective in using smaller LLMs to implement active merchant NPCs. Additionally, we found three irregular cases arising from the responses of LLMs.
footnotetext: Equal contribution.**footnotetext: Corresponding author
1 Introduction
@@
Exchanging in-game items is an integral part of open-world role-playing games because the utility of an item depends on the current attributes of a player. For example, a player with high agility may not benefit from an item that enhances agility. Thus, game developers implement an exchange system in their games to enhance gameplay and item utility.
@@
While games commonly feature merchant non-player characters (NPCs) to facilitate item exchanges, these interactions are typically scripted; the merchant only enables players to buy and sell items at fixed prices, but without the dynamics of real-world transactions. Typically, developers do not allow merchants to alter prices or communicate in free form; instead, they simply present a fixed price to the player without any negotiation. However, recent advancements in LLM-integrated games show growing interest in dynamic and flexible interactions, similar to real-world communication where players and NPCs engage in co-constructed dialogue ReLU Games and KRAFTON (2024); Latitude (2019); Li et al. (2024). To mirror real-world interactions and enhance player immersion, we propose a novel framework that leverages LLMs to enable merchant NPCs to engage in actual negotiations with players.
@@
To create an active merchant NPC, which is more aligned to the real world, we need to address two issues that cause the current merchant to act passively: pricing and communication. First, regarding passive pricing, the merchant has no authority to adjust item prices. Instead, the game developers decide prices based on the items’ utility in the game. However, in the real world, sellers can adjust the prices of their goods according to item specifications. To the best of our knowledge, gaming industry researchers have not adequately filled this gap. Since item descriptions often convey sufficient information about utility, we argue that they can be used to estimate the value of unseen items by leveraging information and values from other known items. Specifically, we let LLMs appraise game items by observing other items.
@@
Second, regarding passive communication, merchants can only interact with players in a scripted manner. Researchers have explored the adoption of LLMs within games to facilitate more immersive player experiences Phillips et al. (2024); Peng et al. (2024). However, the way merchants communicate has been given limited focus, which remains a one-way interaction. Current merchant NPCs simply display a predefined list of items for players to choose from, and players respond by clicking on items to make a purchase. This purchasing experience is uniform across all interactions with a merchant, regardless of the individual player involved in the transaction. To enable two-way communication, we propose a negotiation style that mirrors real-world transactions. Through this study, we explore whether LLMs can be used to foster a negotiating interaction within the merchant context.
@@
In this paper, we propose a framework for developing a More Active meRchanT NPC, called mart. This framework consists of two main components: appraiser and negotiator. The appraiser module addresses the issue of passive pricing by estimating the value of given items. The negotiator module addresses the passive communication issue by negotiating with players. As game developers have diverse requirements when deploying their games, we conducted experiments to compare potential candidates for each module using a public WoW Classic game item dataset. Specifically, we tested both finetuning methods and $n$-shot prompting methods on Llama 3 models, which come with a wide range of parameter sizes up to 405 billion. Our GitHub repository [^1] http://github.com/elu-lab/mart provides implementation details, prompts, and model outputs.
This study has the following contributions:
@@
  • Through statistical and qualitative analyses, we present multiple implementation options for active merchant NPCs, tailored to suit different user preferences.
2 Related Work
2.1 Using LLMs to communicate with NPCs
Recently, game developers have started using LLMs in their games to provide more human-like interaction between players and NPCs. Specifically, there is a growing concern about making conversational NPCs with LLMs Viggiato and Bezemer (2024); Gallotta et al. (2024); Christiansen et al. (2024); Cox and Ooi (2024); Gao et al. (2023); Marincioni et al. (2024). For example, Christiansen et al. used LLM-based NPCs to support player interactions within a murder mystery game by helping players interrogate, collect clues, and explore. Buongiorno et al. introduced a framework for integrating LLM-based NPCs with memory to ensure narrative consistency during free-form player interactions, demonstrated in a turn-based, role-playing detective thriller game.
Despite their contributions, prior studies have largely treated LLM-based NPCs as information providers; That is, they only supported simple interactions where NPCs present knowledge or reasoning to the player. However, more complex forms of interaction that involve mutual influence, such as negotiation, have been largely underexplored in the context of LLM-integrated games.
In real-world applications, NLP researchers have investigated the potential of LLMs to negotiate Jin et al. (2024); Shea et al. (2024); Hua et al. (2024). For example, Jin et al. leveraged LLMs to generate persuasive dialogues based on everyday scenarios and showed that their approach was more persuasive than other models. Similarly, Shea et al. suggested a personal negotiation coach based on LLMs and proved their effectiveness in assisting human negotiators. Inspired by prior work demonstrating the negotiation potential of LLMs, we examine negotiation interactions of an LLM-based NPC in a gaming context.
2.2 Using LLMs to estimate values
As item prices are usually fixed in a game system, research has not sufficiently focused on predicting item prices based on item specifications. An active merchant, however, should have the ability to decide item prices in its shop, similar to a real-world situation. Thus, we designed an appraiser module, inspired by studies on estimating prices of real assets.
In real-world applications, researchers have investigated machine learning models for estimating commodity prices. Early studies attempted to predict prices based on a fixed set of features Patel et al. (2015); Mohamed et al. (2022). For example, Mohamed et al. (2022) used a fixed set of features and a machine-learning method for predicting the price of seasonal goods. However, such models generally do not perform well when predicting prices of out-of-domain goods or newly introduced features. To handle such unseen goods or features, researchers utilized latent representations from textual descriptions of items with a language model Ni et al. (2024); Geng et al. (2024). For instance, Ni et al. demonstrated that LLMs can predict stock market movements from the earnings report of a company to support information-based investment. Extending prior works, we demonstrate that LLMs can also achieve strong predictive performance within games.
3 MART Framework
@@
We propose mart, an LLM-powered framework for active merchant NPC. This framework consists of two main modules: negotiator and appraiser. Inspired by the behavior of real-world merchants, MART autonomously sets prices for its items and engages in price negotiation with players. When a player expresses interest in buying an item, the appraiser module suggests a retail price for the item. Based on this suggestion, the negotiator module begins negotiations with players who wish to buy the item from the merchant NPC. This scenario is different from the purchasing experience of players in current games, as shown in Figure 1.
@@
First, the appraiser module is designed using a language model that interprets an item description to estimate its retail price. The module takes item descriptions presented as natural language sentences, which include details such as required level, effects, and durability. Since the structure of such descriptions can vary across different games, we adopt a natural language format to ensure versatility. In Section 4, we introduce LLM-based appraiser modules using two straightforward approaches and evaluate their performance.
@@
Second, the negotiator module employs a language model to facilitate negotiation dialogues with players. This module receives the item descriptions, retail price, and history of previous conversations. It then generates the appropriate response aimed at persuading the player. Since effective negotiation involves various tactics to persuade players, we experimented LLM-based negotiator modules with two simple approaches that integrate 10 tactics, as detailed in Section 5.
4 Appraiser Module
@@
To develop the appraiser module, we experimented with two approaches: $n$-shot in-context learning (ICL) Dong et al. (2024) and supervised finetuning (SFT). Here, we set the $n$ as 10 for our ICL approach and called LLMs via public APIs [^2] http://openrouter.ai. In this section, we outline each approach and illustrate our findings through an experiment on a WoW Classic item dataset.
4.1 Two tested approaches
@@
First, to leverage the ability of LLMs to generalize from a set of given examples, we tested ICL method. ICL enhances reasoning through demonstrations and requires no additional training, making it a simple and efficient method to implement. Thus, we directly employed LLMs and provided ten random examples in the prompt. We used Llama models, a popular open-source family of LLMs: Llama-3.1 8b, 405b, and 3.2 1b. Note that ICL does not require any additional training to update LLM parameters.
@@
Second, owing to the substantial computational resources and time required by the 405 billion parameter model, we tested the SFT method on smaller LLMs. To avoid catastrophic forgetting, we froze their parameters and trained an additional adapter. When we input an item description, the smaller LLM transforms the description into a single latent vector. Thereafter, the additional adapter uses the vector to predict the retail price of the item as a regression problem. We used Llama 3.1 8b and 3.2 1b.
4.2 Used Dataset
@@
We used an item dataset from WoW (World of Warcraft) Classic  [^3] https://www.kaggle.com/datasets/mylesoneill/classic-world-of-warcraft-auction-data published by Blizzard Entertainment. By extending the existing WoW Classic item dataset, we created a dataset by crawling item information from the provided URLs for each item and retaining only purchasable items with disclosed prices. Dataset construction followed three steps. First, we collected 3,376 items by excluding item derivatives from the crawled dataset to simplify the problem. We then removed 106 items priced below 10 coppers (the cheapest currency unit in WoW) to make enough room for price negotiation during transactions. Finally, we converted all retail prices into coppers, using the conversions of 1 gold to 100 silver and 1 silver to 100 coppers. As a result, we collected a total of 3,270 items, with prices ranging from 10 coppers to 57,018 coppers. More than 50% of the items were priced below 1,250 coppers, with a median price of 1,238 coppers and an average price of 3,249 coppers. We further divided the dataset into a training set (80%, 2,616 items), validation set (10%, 327 items), and test set (10%, 327 items).
4.3 Evaluation metrics
@@
We used four metrics to evaluate the suitability of the two approaches for implementing the appraiser module: mean absolute percentage error (MAPE), standard deviation, skewness, and unexpected output rate (UOR). First, we used MAPE. Due to the wide range of item prices (from 10 to 57,018 coppers), the absolute errors of higher-priced items may overshadow those of lower-priced items. Therefore, we used percentage errors instead of absolute errors. We computed MAPE as the ratio of error to true price, as illustrated in Equation 2, where $\hat{y}_{i}$ is the predicted price and $y^{*}_{i}$ is the true price for an item $i$.
@@
4.4 Results
MAPE Std. Dev. Skewness UOR
ICL 1b 14.68 44.42 10.23 29.50
8b 4.34 12.36 6.85 20.49
@@ 4.4 Results
SFT 1b 3.59 11.57 -5.84 -
8b 2.66 11.26 -3.47 -
Table 1: Results of assessing two methods of appraiser
@@
Table 1 shows the result of our experiments, comparing ICL and SFT methods. Of the five models, the ICL-405b achieved the best performance. It successfully outputted a retail price (94.8% of the cases) with the lowest MAPE (1.34%) and produced a few unexpected outputs (5.2%), while it was slightly skewed toward underestimation (-5.06). The second-best model was the SFT-8b. It exhibited a slightly higher MAPE score (2.66%) and the lowest absolute skewness toward underestimation (-3.47). Meanwhile, the SFT-1b demonstrated lower performance than the best model (3.59%) with a similar underestimation (-5.84). These results are different from those of smaller ICL models. The ICL-1b exhibited very high MAPE (14.68%) with highly overestimated prices (10.23) and a high unexpected output rate (29.05%). Similarly, the ICL-8b overestimated prices (6.85) with a moderate level of MAPE (4.34) and a high unexpected output rate (20.59%).
4.5 Discussion
@@
We discuss our results in terms of three factors–performance, efficiency, and reliability–to assess the in-game applicability of the appraiser module. First, the appraiser has the potential to be used in a game because of its performance. Except for ICL-1b, five models demonstrated a MAPE of less than 5%. While the difference may seem substantial for high-priced items, it does not exceed 100 coppers for over half of the items, considering that the median price was 1,238 coppers. As gold coins are more frequently used in WoW Classic, a difference of 100 coppers is acceptable. Moreover, when using larger ICL models, the appraised prices aligned more closely with the true prices. Therefore, for game developers seeking a more precise appraiser module, ICL-405b is the best option, as it exhibited an appraisal error of less than 16 coppers for the median price.
@@
Second, the SFT method is efficient for developing an appraiser module. Results show that SFT-8b performed much closer to ICL-405b than to ICL-8b; even the SFT-8b and ICL-8b used almost an identical number of parameters. Also, the SFT-1b outperformed the two ICL models, ICL-1b and ICL-8b. These results imply that it is possible to use a smaller LLM to implement an appraiser within a low-resource environment by finetuning it. Note that the ICL models achieved high performance by observing ten random pairs of item descriptions and prices. Researchers have reported that the performance of ICL methods can be improved by carefully curating examples or increasing the number of examples Zhang et al. (2022). However, this curation demands additional human resources, and increasing the number of examples incurs higher computational costs. By contrast, a dataset for SFT can be generated with significantly fewer resources, using pairs of items and their prices.
@@
Third, it is worth noticing that ICL methods can sometimes be unreliable when playing the role of an appraiser. We observed cases in which ICL models produced unexpected outputs. For instance, ICL-405b generated multiple price candidates or a continuous range of prices, despite our request for a single price output. Moreover, smaller ICL models occasionally failed to predict prices. Although we did not adopt any post-processing methods for these outputs, game developers should prevent such failures in their games. Alternatively, developers can use the SFT method, ensuring an LLM predicts a specific price.
5 Negotiator Module
@@
We introduce a negotiator module inspired by real-world merchants, aiming to sell items while pursuing profit. To develop negotiators, we tested two approaches: zero-shot prompting (ZSP) and knowledge distillation (KD). We chose different methods from those used in the appraiser module, as supervised negotiation data is difficult to obtain. In the following subsections, we describe our methods, dataset generation procedure, and details of the experiments. Then, we discuss our findings by comparing these two approaches.
5.1 Two tested approaches